OpenStudy (anonymous):

What is the slope of the line passing through the points (1, –5) and (4, 1)? 2 -4/5 5/4 -2

OpenStudy (alexandervonhumboldt2):

use lope formula: \[\huge \frac{ y_1-y_2 }{ x_1-x_2 }\]

OpenStudy (alexandervonhumboldt2):

*slope

OpenStudy (susanna):

\[\frac{ 1-(-5) }{ 4-1 }=\frac{ 1+5 }{ 3 }=\frac{ 6 }{ 3 }=2\]